OpenStudy (straighta11):

Which of the following inferences can be drawn about the buying pattern of the online customers versus that of the walk-in customers? Online customers buy an average of five packs compared with walk-in customers, who buy an average of three packs. Online customers buy an average of five packs compared with walk-in customers, who buy an average of two packs. Online customers buy an average of six packs compared with walk-in customers, who buy an average of three packs. Online customers buy an average of six packs compared with walk-in customers, who buy an average of two packs. MEDAL

OpenStudy (straighta11):

@ganeshie8 @Kainui

OpenStudy (wolfman77):

Okay, Count the total number of packs bought by on-line customers and the same for walk-in. You can then find the average number of packs per customer by dividing the above totals by the number of customers Do you understand?

OpenStudy (wolfman77):

So, Number packs for on-line customers = 5*6 + 4*5 + 4 = ?

OpenStudy (wolfman77):

5*6 + 4*5 + 4 = 54... So, divide that by 10 to get the mean for online customers. And then what do you have?

OpenStudy (wolfman77):

The answer is: Online customers buy an average of five packs compared with walk-in customers, who buy an average of three packs.

OpenStudy (straighta11):

thank you so much! @Wolfman77 could you help me with another?